OpenStudy (anonymous):

Can someone walk me through Direct Linear Variation? (8th grade)

OpenStudy (anonymous):

What specifically do you need to know?

OpenStudy (anonymous):

Well, my online math curriculum likes to make this really complicated so like this problem: Find the constant of variation for the relationship f(x)= 40x. A.40 B.x C.20 D.f(x)

OpenStudy (anonymous):

A really quick overview would be the following: Direct linear variation is a result of the variable being multiplied or divided by a constant. Addition or subtraction will not work. For example: y=2x where x is the variable and 2 is the constant. When drawn on a graph, it will be a straight line and will pass through point (0,0) In this case the constant will be 40. This is because, no matter what you change x to, y will always be a multiple of 40, hence it is a constant

OpenStudy (anonymous):

Examples of direct linear variation would be: y=2x, f(x)=20x, f(y)=-3y where the number (including the negative) is the constant Examples of non direct linear variation would be: y=9/x, y=x^2, y=2+x

OpenStudy (anonymous):

Ahhh!!! Thanks it makes sense now! Fan and medal. :D